OpenStudy (iotero81):

A sporting goods store sells 3 fishing reels and 4 fishing Rods for $523, later they sell 4 fishing reels and 3 fishing rods for $507 =. Find the price of each item

rebeccaxhawaii (rebeccaxhawaii):

\(\Huge\boxed{\boxed{\boxed{\huge\mathbb{WELCOME~TO~OPENSTUDY} }}}\) Hey i have made this to hopefully answer any questions you may have. But if it doesn't feel free to pm me c: http://openstudy.com/users/rebeccaxhawaii#/updates/56f5205ce4b07a8c82287f75 Good Luck with you studies! \(\Huge\heartsuit\)

rebeccaxhawaii (rebeccaxhawaii):

okay so we will write it like this 3x+4y=523 4x+3y=507 to solve for x and y - got it?

ILovePuppiesLol (ilovepuppieslol):

http://openstudy.com/study#/groups/Mathematics Hello There! Welcome to Openstudy!! Since you are new here, read this tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/56f5205ce4b07a8c82287f75 This is in the wrong section and you would likely to get more help in the Math section! http://openstudy.com/study#/groups/Mathematics

likeabossssssss (likeabossssssss):

plz close this and ask the question in the right section thx